OpenStudy (anonymous):

wildlife biologist catch, tag and release 32 bears at a game reserve. Later 10 bears are caught and 4 of them have tags. Estimate the total amount of bears at the game reserve. a. 320 b. 316 c. 128 d. 80

OpenStudy (anonymous):

help please :(

OpenStudy (anonymous):

you multiply

OpenStudy (anonymous):

what am i suppose to multiply?

OpenStudy (anonymous):

well cross multiply 4/10 = 32/x

OpenStudy (anonymous):

im not getting it, im not that good at math :/

OpenStudy (anonymous):

32*10 divided by 4=?

OpenStudy (anonymous):

80?

OpenStudy (anonymous):

 cross multiply and then you get 32*10 = 4x